# Dexterous Manipulation through Imitation Learning: A Survey

### v4

@@ IV-D Tactile Sensors on Dexterous End-effectors
Tactile sensing is important for enhancing the perception and control capabilities of robotic end-effectors, particularly in dexterous manipulation tasks that require fine contact interaction, compliance, and force regulation. Although previous end-effectors such as two-finger grippers and parallel jaw mechanisms often rely solely on vision or position feedback, modern multi-fingered dexterous hands [249, 250] are increasingly incorporating high-resolution tactile sensors to achieve more stable and versatile manipulation in unstructured or dynamic environments.
In general, tactile sensors provide direct measurements of contact force, pressure distribution, slip, surface texture, and sometimes even temperature or material properties. They are typically mounted on the fingertips, phalanges, or palm surfaces of robotic hands, allowing feedback-rich control policies such as force modulation, slip prevention, and adaptive grasping. In recent years, tactile sensors have become essential for learning-based approaches, where tactile data are used for contact state estimation, affordance recognition, and policy refinement [251, 252, 253, 254].
@@
  • Resistive and Capacitive Tactile Sensors: These are thin, flexible, and cost-effective, and detect pressure changes through variations in electrical resistance or capacitance. They are widely used in fingertip arrays and robotic skin [255, 256, 257].
@@
  • Piezoelectric Sensors: These generate voltage in response to applied stress and are suitable for dynamic contact detection, such as impact and slip sensing [258, 259].
@@
  • Optical and Vision-Based Tactile Sensors: Examples such as GelSight [260], TacTip [261], and GelStereo tip [262] use internal cameras and deformable gel surfaces to reconstruct high-resolution 3D contact geometry. These sensors offer rich, interpretable data and are increasingly used in learning-based manipulation [263].
@@
  • Magnetic and Hall Effect Sensors: Often embedded within soft materials or joints, these detect local deformations or forces via magnetic field changes [264, 265, 266, 267].
@@
  • Bio-inspired Sensors: Inspired by human skin or mechanoreceptors, these sensors aim to capture multi-modal tactile features (e.g., pressure + vibration + shear) in compact designs. Examples include BioTac [268], NeuTouch [269], and GTac [270].
@@
The integration of tactile sensing into robotic hands offers significant advantages. First, it enables robust manipulation under conditions of visual occlusion or low illumination by relying solely on contact feedback. Second, tactile sensors facilitate adaptive grasping strategies such as regrasping, force-limited lifting, and slip compensation. Third, tactile data contribute to improved grasp quality estimation, object property inference, and planning for in-hand manipulation. However, tactile sensing also faces several challenges. Calibration and drift over time, especially in soft or flexible sensors, can reduce measurement fidelity. High-resolution sensors (e.g. GelSight [260]) are relatively bulky and have limited durability under repeated contact. Furthermore, real-time integration of tactile data into control pipelines demands efficient signal processing, accurate contact modeling, and often machine learning-based interpretation.
@@
In the context of IL for dexterous manipulation, tactile sensing provides a rich stream of contact information that complements visual and proprioceptive cues. Demonstrations collected from expert teleoperation or kinesthetic teaching can include synchronized tactile signals, allowing the learning agent to capture subtle contact events, such as incipient slip, rolling, or micro-adjustments of finger force, that are often invisible to vision alone. By incorporating tactile feedback into the policy representation, IL systems can better generalize to novel objects and adapt to variations in shape, compliance, or surface texture. Recent work  [252, 271, 272, 273] showed that tactile-informed IL can substantially improve grasp stability, reduce failure rates under visual occlusion, and enable fine in-hand reorientation without explicit object models.
@@ IV-E1 Morphology:
Redundant Degrees of Freedom and Ergonomic Alignment. High-DoF hands (e.g., Shadow Hand, 20+ DoF) can replicate fine human manipulations such as fingertip control, but they increase the dimensionality of the action space, leading to severe sparsity issues in IL. The RH20T dataset [139] shows that the demonstration data requirement for high-DoF hands grows exponentially. In contrast, three-fingered grippers (e.g., DoraHand [235]) trade off some dexterity for reduced policy learning complexity, achieving better zero-shot generalization in cross-domain tasks such as those in BridgeData V2 [274]. Anthropomorphic hands (e.g., ILDA Hand [148]) with geometric consistency to the human demonstrator’s workspace reduce cross-domain mapping errors. Antotsiou et al. [23] highlighted that differences in morphology, degrees of freedom, and joint constraints between the human hand and the robot hand can amplify retargeting errors and consequently raise the likelihood of task failure.
IV-E2 Actuation:
@@
Trade-Offs Between Transmission Dynamics and Controllability. Hands such as the Shadow Hand use tendon-driven actuation, where compliance introduces nonlinear friction and hysteresis, increasing noise in the action–state mapping of demonstration data. Tendon-driven systems introduce nonlinearities such as friction and hysteresis in the transmission chain, which are known to degrade control fidelity in dexterous manipulators [187]. Rigid linkage actuation (e.g., BarrettHand [222]) reduces dynamic errors, but underactuated designs (e.g., adaptive grasping) limit the reproduction of fine finger motions. Direct-drive hands (e.g., LEAP Hand [209]) achieve millimeter-level control precision through joint-level motors but suffer from increased inertia due to motor mass, requiring IL policies to learn dynamic feedforward compensation.
IV-E3 Sensor Configuration:
@@
Complementarity in Multimodal Perception. High-resolution tactile sensing (e.g., GelSight  [260]) can capture micro-force adjustments and slip signals in human demonstrations, significantly improving IL’s understanding of contact dynamics. Huang et al. [252] reported that removing tactile input causes task success rates to plummet to near-random levels, underscoring the critical importance of tactile information for successful multi-object in-hand manipulation. Differences in sampling frequency (e.g., vision at 30 Hz vs. tactile sensing at 1 kHz) introduce synchronization challenges. Lin et al. [285] addressed this via a cross-modal alignment network, reducing motion consistency error by 25% in bimanual coordination tasks. NeuralFeels [275] aligns vision with tactile streams via global timestamps and fuses both modalities in a joint neural implicit field, cutting trajectory-consistency error by 25% for precise in-hand manipulation.
@@
Future end-effector designs should follow a “task–morphology–algorithm” co-optimization paradigm. High-dexterity tasks (e.g., surgical suturing) favor fully actuated hands with dense tactile arrays (e.g., Shadow Dexterous Hand [150]) and employ hierarchical IL frameworks [74] to mitigate the curse of dimensionality. Low-cost generalization tasks (e.g., household tidying) use three-fingered grippers with vision-dominant sensing, combined with domain randomization [276] to improve cross-object robustness. Real-time collaboration tasks (e.g., human–robot co-manipulation) balance actuation latency and control precision. For instance, RAPID Hand [277] achieves <7<7 ms response delay through lightweight design.
TABLE VI: Comparison of Teleoperation Data Collection Approaches
@@
Vision-Based Systems [278, 18, 17, 78, 279, 20, 19, 280, 114, 281] Non-contact image-based motion capture Single or multi-camera configurations RGB or RGB-D inputs, frequently accompanied by hand pose estimation Low-cost hardware (especially single camera) Easy to deploy and user-friendly No wearables required Susceptible to occlusion and lighting conditions Limited 3D precision and depth estimation Latency due to vision processing
Mocap Gloves [282, 16, 283] Wearable gloves with IMUs, flex sensors, or magnetic sensors Real-time hand joint tracking Often used with external cameras or LiDAR High joint-level accuracy (millimeter-level) Fast response time Robust under varied visual conditions High equipment cost Requires calibration and setup Potential discomfort for extended use
VR/AR Controllers [14, 284, 285, 283, 286, 287, 288] Head-mounted display with spatial tracking Hand-held controllers with 6-DoF input Immersive virtual environments Intuitive and interactive control Multimodal feedback (e.g., visual and occasional haptic) Consumer-grade accessibility Dependence on virtual environments Challenging deployment and non-negligible control latency Limited haptic realism and force feedback
Exoskeleton and Bilateral Systems [289, 290, 291, 292, 77, 143, 293, 144] Direct joint-space mapping Includes bilateral force and torque feedback Leader-follower robot configurations High-fidelity control suitable for precision tasks Enables tactile interaction and real-time feedback Complex and bulky setups Expensive hardware and maintenance Restricted operator mobility
@@ V Dexterous Manipulation with Teleoperation and Video Demonstration
This section discusses two main approaches to dexterous manipulation: acquiring demonstration data through teleoperation, and learning directly from demonstration videos. Teleoperation facilitates direct human control over robotic systems, leveraging human expertise for learning complex manipulation tasks, while video-based learning harnesses rich visual data to enable autonomous skill acquisition from natural human demonstrations. Together, these approaches address challenges in teaching robots fine-grained manipulation behaviors across diverse scenarios. The section further discusses datasets and benchmarks in supporting imitation learning frameworks that underpin these methodologies.
@@ V-A Teleoperation Systems for Dexterous Manipulation
A typical teleoperation system consists of two main components: the local site and the remote site, as demonstrated in Fig. 5. The local site includes a human operator and a suite of interactive I/O(input/output) devices. The output devices provide real-time status about the remote robot and its surrounding environment, while the input devices allow the operator to issue commands in diverse forms, thereby controlling the remote robot’s actions. The remote site primarily contains the robot itself, which is equipped with various sensors to gather perceptions of its state and the surrounding environment. Upon receiving teleoperation commands from humans, the robots can perform the corresponding actions and complete tasks.
Figure 5: Teleoperation Frameworks and Commonly Used Devices. (a) mocap gloves (b) VR controllers (c) joystick (d) camera (e) exoskeleton (f) dexterous hand (g) dual-arm robot (h) robot arm.
To accurately convey human operators’ intentions to robotic systems, previous works have employed a wide range of human-robot interaction devices. Human operators with work experience can easily identify the current state of a robot through the image; however, accurately translating human instructions to robot actions remains a challenge. Some traditional controllers act on this: 1) joysticks [294] 2) haptic devices [295]; However, manipulation tasks often involve delicate movements and complex interactions, such as grasping, moving, and positioning small or irregularly shaped objects. These tasks necessitate devices that can offer dexterous interfaces to ensure the safety and efficacy of the robot’s actions. Precision and real-time feedback are crucial. Commonly used devices include: 1) cameras [18, 296, 17, 19, 280, 278, 279, 78, 20]; 2) mocap gloves [297, 298, 282, 299, 300, 301]; 3) VR/AR controllers [302, 303, 14, 304, 305, 78, 284, 27, 286, 306, 307, 308]; 4) exoskeletons and bilateral systems [289, 291, 292, 77, 143, 293]. A comparative summary of representative teleoperation approaches is presented in Tab. VI, highlighting their key characteristics, advantages, and limitations to facilitate effective human-robot interaction.
@@ V-A1 Vision-based Teleoperation Systems
Recently, advancements in computer vision have led to the development of vision-based teleoperation systems. However, their accuracy in capturing hand movements is often compromised by factors such as occlusion, lighting, resolution, background, and inaccurate 3D estimation issues. Several methods have been proposed for robust hand pose estimation and reliable mapping to the robot end-effector. Li et al. [278] developed a vision-based teleoperation system by training TeachNet on pairs of images of human hands and simulated robots to form mappings between a human hand and a robotic Shadow Hand in the latent space. Dexpilot [18] utilized a calibrated multi-camera system to estimate hand poses to teleoperate an Allegro Hand. Riemannian Motion Policies (RMPs) are employed to compute the Cartesian pose of the hand, facilitating hand-arm motion control. Subsequent approaches, such as Robotic Telekinesis [17] and DIME [78], simplified requirements to a single RGB camera, thus reducing the need for calibration. This is achieved through a general mapping method between humans and robots that have different kinematic structures. Additionally, Robotic telekinesis [17] adjusts the position and orientation of the end-effector relative to its base using the relative position and direction of the human wrist to the torso, enabling the teleoperate both arm and hand. However, these methods still suffer from occlusion issues due to the single fixed camera setting. To solve this problem, Transteleop [279] introduced a system that utilizes real-time active vision with a depth camera mounted on the end-effector of the remote UR5 robot arm. During teleoperation, this robot arm can reposition the camera to enhance its field of view and improve hand pose estimation accuracy.
The morphology discrepancy between the human hand and the robot hand might impede the operator from intuitively controlling the robot. To address this, Qin et at. [20] developed a user-friendly interface by constructing a customized robotic hand modeled after the specific shape of a human hand. Demonstrations performed with this customized robot hand can be directly transferred to any dexterous robot hand. AnyTeleop [19] proposed a solution to the self-occlusion problem by integrating images from multiple cameras, each offering different perspectives. To further enhance precision in observation, ACE [280] mounted the camera under the end-effector of the exoskeleton to maintain a clear view of the hands and wrists. MimicPlay [114] employs two calibrated cameras in different viewpoints to reconstruct 3D hand locations. The teleoperation data for the robot is collected using the RoboTurk system [281], which operates via an IMU-equipped smartphone.

### v5

@@ IV-D Tactile Sensors on Dexterous End-effectors
Tactile sensing is important for enhancing the perception and control capabilities of robotic end-effectors, particularly in dexterous manipulation tasks that require fine contact interaction, compliance, and force regulation. Although previous end-effectors such as two-finger grippers and parallel jaw mechanisms often rely solely on vision or position feedback, modern multi-fingered dexterous hands [ueno2024multi, lin2025pp] are increasingly incorporating high-resolution tactile sensors to achieve more stable and versatile manipulation in unstructured or dynamic environments.
In general, tactile sensors provide direct measurements of contact force, pressure distribution, slip, surface texture, and sometimes even temperature or material properties. They are typically mounted on the fingertips, phalanges, or palm surfaces of robotic hands, allowing feedback-rich control policies such as force modulation, slip prevention, and adaptive grasping. In recent years, tactile sensors have become essential for learning-based approaches, where tactile data are used for contact state estimation, affordance recognition, and policy refinement [lee2024dextouch, huang2025dih, heng2025vitacformer, zhang2025kinedex].
@@
  • Resistive and Capacitive Tactile Sensors: These are thin, flexible, and cost-effective, and detect pressure changes through variations in electrical resistance or capacitance. They are widely used in fingertip arrays and robotic skin [hailiang2024flexible, liu2024material, chen2025capacitive].
@@
  • Piezoelectric Sensors: These generate voltage in response to applied stress and are suitable for dynamic contact detection, such as impact and slip sensing [wang2025piezotac, tang2025recent].
@@
  • Optical and Vision-Based Tactile Sensors: Examples such as GelSight [yuan2017gelsight], TacTip [ward2018tactip], and GelStereo tip [zhang2025gelstereo] use internal cameras and deformable gel surfaces to reconstruct high-resolution 3D contact geometry. These sensors offer rich, interpretable data and are increasingly used in learning-based manipulation [sun2025tactile].
@@
  • Magnetic and Hall Effect Sensors: Often embedded within soft materials or joints, these detect local deformations or forces via magnetic field changes [yan2021soft, yan2022tactile, park2023magtac, ding2024soft].
@@
  • Bio-inspired Sensors: Inspired by human skin or mechanoreceptors, these sensors aim to capture multi-modal tactile features (e.g., pressure + vibration + shear) in compact designs. Examples include BioTac [wettels2014multimodal], NeuTouch [taunyazov2020event], and GTac [lu2022gtac].
@@
The integration of tactile sensing into robotic hands offers significant advantages. First, it enables robust manipulation under conditions of visual occlusion or low illumination by relying solely on contact feedback. Second, tactile sensors facilitate adaptive grasping strategies such as regrasping, force-limited lifting, and slip compensation. Third, tactile data contribute to improved grasp quality estimation, object property inference, and planning for in-hand manipulation. However, tactile sensing also faces several challenges. Calibration and drift over time, especially in soft or flexible sensors, can reduce measurement fidelity. High-resolution sensors (e.g. GelSight [yuan2017gelsight]) are relatively bulky and have limited durability under repeated contact. Furthermore, real-time integration of tactile data into control pipelines demands efficient signal processing, accurate contact modeling, and often machine learning-based interpretation.
@@
In the context of IL for dexterous manipulation, tactile sensing provides a rich stream of contact information that complements visual and proprioceptive cues. Demonstrations collected from expert teleoperation or kinesthetic teaching can include synchronized tactile signals, allowing the learning agent to capture subtle contact events, such as incipient slip, rolling, or micro-adjustments of finger force, that are often invisible to vision alone. By incorporating tactile feedback into the policy representation, IL systems can better generalize to novel objects and adapt to variations in shape, compliance, or surface texture. Recent work  [huang2025dih, ablett2024multimodal, guzey2024see, murooka2025tact] showed that tactile-informed IL can substantially improve grasp stability, reduce failure rates under visual occlusion, and enable fine in-hand reorientation without explicit object models.
@@ IV-E1 Morphology
Redundant Degrees of Freedom and Ergonomic Alignment. High-DoF hands (e.g., Shadow Hand, 20+ DoF) can replicate fine human manipulations such as fingertip control, but they increase the dimensionality of the action space, leading to severe sparsity issues in IL. The RH20T dataset [fang_rh20t_2024] showed that the demonstration data requirement for high-DoF hands grows exponentially. In contrast, three-fingered grippers (e.g., DoraHand [wang_dorahand_2022]) traded off some dexterity for reduced policy learning complexity, achieving better zero-shot generalization in cross-domain tasks such as those in BridgeData V2 [walke2023bridgedata]. Anthropomorphic hands (e.g., ILDA Hand [kim_integrated_2021]) with geometric consistency to the human demonstrator’s workspace reduced cross-domain mapping errors. Antotsiou et al. [antotsiou2018task] highlighted that differences in morphology, degrees of freedom, and joint constraints between the human hand and the robot hand can amplify retargeting errors and consequently raise the likelihood of task failure.
IV-E2 Actuation
@@
Trade-Offs Between Transmission Dynamics and Controllability. Hands such as the Shadow Hand use tendon-driven actuation, where compliance introduces nonlinear friction and hysteresis, increasing noise in the action–state mapping of demonstration data. Tendon-driven systems introduce nonlinearities such as friction and hysteresis in the transmission chain, which degrades control fidelity in dexterous manipulators [grebenstein_approaching_2014]. Rigid linkage actuation (e.g. BarrettHand [BarrettHand_web]) reduces dynamic errors, but underactuated designs (e.g. adaptive grasping) limit the reproduction of fine finger motions. Direct-drive hands (e.g. LEAP Hand [shaw_leap_2023]) achieve millimeter-level control precision through joint-level motors but suffer from increased inertia due to motor mass, requiring IL policies to learn dynamic feedforward compensation.
IV-E3 Sensor Configuration
@@
Complementarity in Multimodal Perception. High-resolution tactile sensing (e.g., GelSight  [yuan2017gelsight]) can capture micro-force adjustments and slip signals in human demonstrations, significantly improving IL’s understanding of contact dynamics. Huang et al. [huang2025dih] reported that removing tactile input causes task success rates to plummet to near-random levels, underscoring the critical importance of tactile information for successful multi-object in-hand manipulation. Differences in sampling frequency (e.g., vision at 30 Hz vs. tactile sensing at 1 kHz) introduce synchronization challenges. Lin et al. [285] addressed this via a cross-modal alignment network, reducing motion consistency error by 25% in bimanual coordination tasks. NeuralFeels [suresh2024neuralfeels] aligned vision with tactile streams via global timestamps and fuses both modalities in a joint neural implicit field, cutting trajectory-consistency error by 25% for precise in-hand manipulation.
@@
Future end-effector designs should follow a “task–morphology–algorithm” co-optimization paradigm. High-dexterity tasks (e.g., surgical suturing) favor fully actuated hands with dense tactile arrays (e.g., Shadow Dexterous Hand [shadow_hand_web2]) and employ hierarchical IL frameworks [kim2025srt] to mitigate the curse of dimensionality. Low-cost generalization tasks (e.g., household tidying) use three-fingered grippers with vision-dominant sensing, combined with domain randomization [tobin2017domain] to improve cross-object robustness. Real-time collaboration tasks (e.g., human–robot co-manipulation) balance actuation latency and control precision. For instance, RAPID Hand [wan2025rapid] achieved <7<7 ms response delay through a lightweight design.
TABLE VI: Comparison of Teleoperation-based Data Collection Approaches
@@
Vision-Based Systems [8794277, handa2020dexpilot, sivakumar2022robotic, arunachalam2023dexterous, li2022dexterous, qin2022one, qin2023anyteleop, yang2024ace, wang2023mimicplay, mandlekar2018roboturk] Non-contact image-based motion capture Single or multi-camera configurations RGB or RGB-D inputs, frequently accompanied by hand pose estimation Low-cost hardware (especially single camera) Easy to deploy and user-friendly No wearables required Susceptible to occlusion and lighting conditions Limited 3D precision and depth estimation Latency due to vision processing
Mocap Gloves [caeiro2021systematic, wang2024dex, mosbach2022accelerating] Wearable gloves with IMUs, flex sensors, or magnetic sensors Real-time hand joint tracking Often used with external cameras or LiDAR High joint-level accuracy (millimeter-level) Fast response time Robust under varied visual conditions High equipment cost Requires calibration and setup Potential discomfort for extended use
VR/AR Controllers [zhang2018deep, arunachalam2023holo, radosavovic2023real, mosbach2022accelerating, ding2024bunny, cheng2024opentelevision, lin2024learning] Head-mounted display with spatial tracking Hand-held controllers with 6-DoF input Immersive virtual environments Intuitive and interactive control Multimodal feedback (e.g., visual and occasional haptic) Consumer-grade accessibility Dependence on virtual environments Challenging deployment and non-negligible control latency Limited haptic realism and force feedback
Exoskeleton and Bilateral Systems [falck2019vito, falck2018human, fang2024airexo, kim2023training, zhao2023learning, fu2024mobile, wu2024gello, chi2024universal] Direct joint-space mapping Includes bilateral force and torque feedback Leader-follower robot configurations High-fidelity control suitable for precision tasks Enables tactile interaction and real-time feedback Complex and bulky setups Expensive hardware and maintenance Restricted operator mobility
@@ V Dexterous Manipulation with Teleoperation and Video Demonstration
This section discusses two main approaches to dexterous manipulation: acquiring demonstration data through teleoperation and learning directly from demonstration videos. Teleoperation facilitates direct human control over robotic systems, leveraging human expertise for learning complex manipulation tasks, while video-based learning harnesses rich visual data to enable autonomous skill acquisition from natural human demonstrations. Together, these approaches address challenges in teaching robots fine-grained manipulation behaviors across diverse scenarios. The section further discusses datasets and benchmarks in supporting imitation learning frameworks that underpin these methodologies.
@@ V-A Teleoperation Systems for Dexterous Manipulation
A typical teleoperation system consists of two main components: the local site and the remote site, as demonstrated in Fig. 5. The local site includes a human operator and a suite of interactive I/O devices. The output devices provide real-time status about the remote robot and its surrounding environment, while the input devices allow the operator to issue commands in diverse forms, thereby controlling the remote robot’s actions. The remote site primarily contains the robot itself, which is equipped with various sensors to gather perceptions of its state and the surrounding environment. Upon receiving teleoperation commands from humans, the robots can perform the corresponding actions and complete tasks.
Figure 5: Teleoperation frameworks and commonly used devices: (a) mocap gloves, (b) VR controllers, (c) joystick, (d) RGB-D camera, (e) exoskeleton, (f) dexterous hand, (g) dual-arm robot, (h) single-arm robot.
To accurately convey human operators’ intentions to robotic systems, previous works have employed a wide range of human-robot interaction devices. Human operators with work experience can easily identify the current state of a robot through the image; however, accurately translating human instructions to robot actions remains a challenge. Some traditional controllers act on this: 1) joysticks [1041657] 2) haptic devices [toedtheide2023force]; However, manipulation tasks often involve delicate movements and complex interactions, such as grasping, moving, and positioning small or irregularly shaped objects. These tasks necessitate devices that can offer dexterous interfaces to ensure the safety and efficacy of the robot’s actions. Precision and real-time feedback are crucial. Commonly used devices include: 1) cameras [handa2020dexpilot, song2020grasping, sivakumar2022robotic, qin2023anyteleop, yang2024ace, 8794277, li2022dexterous, arunachalam2023dexterous, qin2022one]; 2) mocap gloves [6696638, 10.1145/3197517.3201399, caeiro2021systematic, kumar2015mujoco, li2020mobile, taheri2020grab]; 3) VR/AR controllers [hedayati2018improving, seo2023deep, zhang2018deep, duan2023ar2d2trainingrobotrobot, jang2022bc, arunachalam2023dexterous, arunachalam2023holo, rajeswaran2018learning, ding2024bunny, 8003431, 8594043, 9340822]; 4) exoskeletons and bilateral systems [falck2019vito, fang2024airexo, kim2023training, zhao2023learning, fu2024mobile, wu2024gello]. A comparative summary of representative teleoperation approaches is presented in Table VI, highlighting their key characteristics, advantages, and limitations to facilitate effective human-robot interaction.
@@ V-A1 Vision-based Teleoperation Systems
Recently, advancements in computer vision have led to the development of vision-based teleoperation systems. However, their accuracy in capturing hand movements is often compromised by factors such as occlusion, lighting, resolution, background, and inaccurate 3D estimation issues. Several methods have been proposed for robust hand pose estimation and reliable mapping to the robot end-effector. Li et al. [8794277] developed a vision-based teleoperation system by training TeachNet on pairs of images of human hands and simulated robots to form mappings between a human hand and a robotic Shadow Hand in the latent space. Dexpilot [handa2020dexpilot] utilized a calibrated multi-camera system to estimate hand poses to teleoperate an Allegro Hand. Riemannian Motion Policies (RMPs) are employed to compute the Cartesian pose of the hand, facilitating hand-arm motion control. Subsequent approaches, such as Robotic Telekinesis [sivakumar2022robotic] and DIME [arunachalam2023dexterous], simplified requirements to a single RGB camera, thus reducing the need for calibration. This is achieved through a general mapping method between humans and robots that have different kinematic structures. Additionally, Robotic telekinesis [sivakumar2022robotic] adjusted the position and orientation of the end-effector relative to its base using the relative position and direction of the human wrist to the torso, enabling the teleoperation of both arm and hand. However, these methods still suffer from occlusion issues due to the single fixed camera setting. To solve this problem, Transteleop [li2022dexterous] introduced a system that utilizes real-time active vision with a depth camera mounted on the end-effector of the remote UR5 robot arm. During teleoperation, this robot arm can reposition the camera to enhance its field of view and improve hand pose estimation accuracy.
The morphology discrepancy between the human hand and the robot hand might impede the operator from intuitively controlling the robot. To address this, Qin et al. [qin2022one] developed a user-friendly interface by constructing a customized robotic hand modeled after the specific shape of a human hand. Demonstrations performed with this customized robot hand can be directly transferred to any dexterous robot hand. AnyTeleop [qin2023anyteleop] proposed a solution to the self-occlusion problem by integrating images from multiple cameras, each offering different perspectives. To further enhance precision in observation, ACE [yang2024ace] mounted the camera under the end-effector of the exoskeleton to maintain a clear view of the hands and wrists. MimicPlay [wang2023mimicplay] employed two calibrated cameras in different viewpoints to reconstruct 3D hand locations. The teleoperation data for the robot was collected using the RoboTurk system [mandlekar2018roboturk], which operated via an IMU-equipped smartphone.